Data was collected for 300 fish in the North Atlantic.
The length of the fish (in mm)is summarized in the GFDT below.
Lengths (mm) Frequency
80-89 1
90-99 16
100-109 71
110-119 108
120-129 83
130-139 18
140-149 3
What is the class boundary between the third and fourth classes?
class boundary=

Answer by CPhill(1959)

The third class is 100-109 mm, and the fourth class is 110-119 mm.
The class boundary between these two classes is the midpoint between the upper limit of the third class and the lower limit of the fourth class.
Class boundary = (109 + 110) / 2 = 219 / 2 = 109.5 mm
So, the class boundary between the third and fourth classes is 109.5 mm.